suppose tht 12 dimes and nickels are worth 95 cents . how many dimes and nickels are there??
Answer by elima(1433)

suppose that 12 dimes and nickels are worth 95 cents . how many dimes and nickels are there??
x = nickels
y= dimes
x+y=12
.05x+.1y=.95
x=12-y
.05(12-y)+.1x=.95
.6-.05y+.1y=.95
.05y=.35
y=7
x+(7)=12
x=12-7
x=5
5 nickels
7 dimes
:)